OpenStudy (anonymous):

Does 7|11! Explain why or why not

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

it's not the same

OpenStudy (anonymous):

I do not understand the question then

OpenStudy (anonymous):

7/11 is a good store. So yes, 7/11! is a true statement.

OpenStudy (anonymous):

is the question whether 11 divides seven factorial?

OpenStudy (anonymous):

i'm not really sure. that's just what the problem looks like

OpenStudy (anonymous):

if so, the answer is yes, because \(11!=11\times 10\times 9\times 8\times 7\times 6\times 5\times 4\times 3\times 2\)

OpenStudy (anonymous):

i.e. \(11!\) is a multiple of \(7\)

OpenStudy (anonymous):

it's not showing the work. it says math processing error

OpenStudy (anonymous):

i do not think so.. \ think it means,\[7\over11!\]

OpenStudy (anonymous):

math proc. error

OpenStudy (anonymous):

@mathsucksalot123 again, go for javascript based ones like Google calculator

OpenStudy (anonymous):

Satellite,please verify?

OpenStudy (anonymous):

if the question is "does" then it must be "does 7 divide eleven factorial

OpenStudy (anonymous):

@help123please. refresh your browser

OpenStudy (anonymous):

Oh right - something must be wrong with OS again.

OpenStudy (anonymous):

i still couldn't see the answer satellit

OpenStudy (anonymous):

so, if what I understand is correct, them the amswer would be a clear no. 7 is not divisible by 11!

OpenStudy (anonymous):

i will write it without the latex 11! = 11*10*9*8*7*6*5*4*3*2

OpenStudy (anonymous):

so 7 divides it for sure

OpenStudy (anonymous):

but if it is like what satellite thinks it is, then it is the same problem as we did before but in different words

OpenStudy (anonymous):

i don't know.I just don't know why he would include them both.